OpenStudy (anonymous):

A jar contained 6 red marbles, 10 green marbles, and 8 blue marbles. A marble was selected at random, the color was recorded, and the marble was placed back in the jar. The table shows the results after 50 trials. What was the relative frequency of selecting a red marble? Round to the nearest percent, if necessary.

OpenStudy (anonymous):

http://prntscr.com/74eckt

OpenStudy (anonymous):

@Butterfield1215

OpenStudy (anonymous):

@butterflydreamer @Vianne

OpenStudy (anonymous):

.

OpenStudy (vianne):

Lightshot doesn't seem to be responding to me. :(

OpenStudy (anonymous):

Ok relative frequency is basically the amount of times that you selected a red marble divided by the total amount of trials

OpenStudy (anonymous):

How many times did you select a red marble? out of the 50 trials Thats just another way of looking at it

OpenStudy (anonymous):

relative frequency is basically the outcomes that occurred but in percentage format

OpenStudy (anonymous):

Relative frequency of selecting red =\(\frac{\text{amt of times selected red}}{\text{ total amt of trials}}\)

OpenStudy (vianne):

Frequency of a particular outcome/total number of experiements.

OpenStudy (vianne):

So total up 10 + 24 + 16 = ?

OpenStudy (anonymous):

50

OpenStudy (vianne):

Ok And there are 10 Reds. Now divide :D 10/50 = ?

OpenStudy (anonymous):

0.2

OpenStudy (anonymous):

20?

OpenStudy (anonymous):

is 20 the answer?

OpenStudy (vianne):

Yep!

OpenStudy (anonymous):

ty

OpenStudy (vianne):

No problem! :)